Please help me solve this word problem. Sat Tranh took a short hike with his friends up Mt. Wachusett. They hiked uphill at a steady pace of 1.5 miles per hour and downhill at a rate of 4 miles per hour. If the time to climb the mountain took an hour more than the time to hike down, how long was the entire hike? This is the exact word problem from the book. Please Help! Thanks. Michael Armani
Found 2 solutions by josgarithmetic, stanbon:
Answer by josgarithmetic(39623)

r*h=d
r is rate in miles per hour
h is time in hours
d is distance in miles

Direction________speed__________time_____________distance
UP_______________1.5____________h+1______________1.5(h+1)=d
DOWN_____________4______________h________________4h=d

The distance up is the same as the distance down.

Answer by stanbon(75887)

Tranh took a short hike with his friends up Mt. Wachusett. They hiked uphill at a steady pace of 1.5 miles per hour and downhill at a rate of 4 miles per hour. If the time to climb the mountain took an hour more than the time to hike down, how long was the entire hike?
-----
Up DATA:
rate = 1.5 mph ; time = x hrs ; distance = 1.5x miles
-----------------
Down DATA:
rate = 4 mph ; time = x-1 hrs ; distance = 4(x-1) miles
-----------------
Equation:
dist up = dist down
1.5x = 4x - 4
2.5x = 4
x = 1.6 hours (time up)
x-1 = 0.6 hours (time down)
-----
Ans: Total time:: 1.6 + 0.6 = 2.2 hours
----------------------
Cheers,
Stan H.